OpenStudy (anonymous):

Wouldn't the measure of side CB be none of the above out of 8, 514, 13, 12, and none of the above?

OpenStudy (anonymous):



OpenStudy (anonymous):

I figured 17 needed to be 153 + x^2, so then x^2= 289-153=36, and the sqrt of 136 is 11.6619037897, so none of the avove.. right?

OpenStudy (anonymous):

some one please help me